if the area of a rectangle is 7200 ft squared and the length is 120 ft what is the length of its diagonal in feet to the nearest whole number?

Answer by ewatrrr(24785)


Hi,
7200 ft squared and the length is 120 ft
7200/120 = 60 = width
Applying the Pythagorean Theorem, with d representing the diagonal
60^2 + 120^2 = d^2
134ft = d, the diagonal in feet to the nearest whole number